jerome's father is three times as old as jerome. Four years ago, he was five times as old. find the age of jerome and his father.

Found 2 solutions by Greenfinch, richwmiller:
Answer by Greenfinch(383)

Say Jerome is now X
Then his father is now 3X
Four years ago, Jerome was X-4
His father was 3X - 4 and was five times older than Jerome
so 3x - 4 = 5(X-4)
3X -4 = 5X - 20
3X + 16 = 5X
16 = 2X
X = 8
To check, Now Jerome is 8, father is 24 yes, three times
Four years ago Jerome was 4, father was 20 yes, five times

Answer by richwmiller(17219)

f=3j
(f-4)=5*(j-4)
substitute 3j for f
3j-4=5*(j-4)
3j-4=5j-20
add 20
3j+16=5j
subtract 3j
16=2j
j=8
f=24